Given the circle X^2+y^2=25 in the first quadrant, find the equation of the tangent line at x=3
Answer by Alan3354(69443)

Given the circle X^2+y^2=25 in the first quadrant, find the equation of the tangent line at x=3
---------
At x = 3, y = 4
----
Slope of the line thru the Origin to (3,4) = 4/3
Tangent line is perpendicular to the radial line.
-----
Slope of tangent line = -x/y = -3/4
-----
y-4 = (-3/4)*(x-3)